OpenStudy (anonymous):

differentiation.... if Y=ln(sinX),then show,d^3Y?dX^3=2cosX/sin^3X ??????????

OpenStudy (anonymous):

third derivative right?

OpenStudy (anonymous):

sorry.....i am gonna write it again

OpenStudy (anonymous):

if y=ln(sinx),then show,d^3y/dx^3y=2cosx/sin^3x????

OpenStudy (anonymous):

dy/dx = cos(x)/sin(x)= cot(x) d^2y/dx^2= -csc^2(x) \[\frac{d^3y}{\text{dx}^3}=2 \text{Cot}(x)\text{Csc}^2(x)\] \[\frac{2 \cos (x)}{\sin (x) }\frac{1}{\sin ^2(x)}\]

OpenStudy (anonymous):

thanks a lotttttttttt

OpenStudy (anonymous):

i've got ittttt